OpenStudy (anonymous):

i need help with this problem Find the particular solution of the differential equation (dy/dx) +4y=8 satisfying the initial condition y(0)= 0

OpenStudy (anonymous):

hi can u help plz

OpenStudy (anonymous):

i got y=-2e^(4x)+2 but its wrong can some one show me where the wrong

OpenStudy (anonymous):

how u get 4xy

OpenStudy (anonymous):

integral

OpenStudy (anonymous):

integral of what

OpenStudy (anonymous):

im a bit rusty, but that should b it.

OpenStudy (anonymous):

its wrong coz i tried it on the web work it say is not correct

OpenStudy (anonymous):

coz its online homework

OpenStudy (anonymous):

dy/dx = 8 - 4y dy / (8-4y) = dx u = 8 - 4y du/dy = -4 du/-4 = dy du / -4u = dx (lnu)/-4 = x + C ln(8-4y) = -4x + C Ce^-4x = 8 - 4y y = 2 - Ce^-4x y(0) = 2 - C = 0 C = 2 y = 2- 2e^-4x verification : dy/dx = 8e^-4x 8e^-4x = 8 - 4y ? 8e^-4x = 8 -4(2-2e^-4x) 8e^-4x = 8e^-4x

OpenStudy (anonymous):

thanks its correct now

OpenStudy (across):

We have that\[\frac{dy}{dx}+4y=8.\]This is a first-order, linear ODE. As such, it can be solved using an integrating factor.\[\mu(x)=e^{\int4dx}=e^{4x}.\]Now you multiply the equation by this integrating factor to obtain\[e^{4x}\frac{dy}{dx}+4e^{4x}y=8e^{4x}.\]The LHS of this equation is the expansion of the product rule\[\frac{d}{dx}[e^{4x}y]=8e^{4x}.\]Integrating, we see\[e^{4x}y=8\cdot\frac{1}{4}e^{4x}+c=2e^{4x}+c.\]Solving for y\[y=2+\frac{c}{e^{4x}}.\]Using the initial value\[0=2+\frac{c}{e^{4\cdot0}},\]\[0=2+c,\]\[c=-2.\]Therefore,\[y=2-\frac{2}{e^{4x}}.\]

OpenStudy (across):

Nevermind; too late. ;P

OpenStudy (anonymous):

ok thats right but can u help me with this one too

OpenStudy (anonymous):

yet much better lol

OpenStudy (turingtest):

this is a good page for these: http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (anonymous):

Solve the separable differential equation (dy/dx)=(-0.7/cosy) and find the particular solution satisfying the initial condition y(0)=pi/3

OpenStudy (anonymous):

thanks that good page i ll red it

OpenStudy (anonymous):

i got sin y=-0.7x +0.866 but i need to solve for y not sin y

OpenStudy (turingtest):

just do inverse sin both sides

OpenStudy (anonymous):

cosydy = -0.7dx siny = -0.7x + C y = arcsin(-0.7x + C) arcsin(C) = pi/3 C = sin(pi/3) C = (3^0.5)/2 y = arcsin(-0.7x + (3^0.5)/2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so jus the arc of it will be fine

OpenStudy (anonymous):

i didnt know that thanks :)

OpenStudy (anonymous):

what about this q: Find the particular solution of the differential equation )x^2dy)/(y^2-3dx)=(1/2y) satisfying the initial condition y(1)=(4)^(1/2)

OpenStudy (turingtest):

are those dy's and dx's in the right place?

OpenStudy (anonymous):

ya

OpenStudy (turingtest):

if it's not supposed to be x^2/(y^2-3)(dy/dx)=1/2y I didn't even know you could write something like that. And why are the conditions written as y(1)=4^(1/2) and not just y(1)=2 ? I think I need to appeal to someone greater than I.

OpenStudy (anonymous):

yes that right

OpenStudy (across):

You crack me up TuringTest, in a good way. :P <3

OpenStudy (turingtest):

tell me across, does the expression we were given look doable?

OpenStudy (across):

I gave it a second look and was perplexed to see a dx thrown in the denominator with no proportional connection to dy whatsoever. He must have typoed it.

OpenStudy (across):

talarowen, make sure you write the problems exactly as they're given if you ever want us to help you solve them. :p